# Are We Truly Forgetting? A Critical Re-examination of Machine Unlearning Evaluation Protocols

### v2

@@ 3.1 Problem Setup and Evaluation Metrics for Unlearning
Logit-based evaluation metrics.    As a classical evaluation protocol, unlearning models are evaluated on the following metrics: Forget Accuracy (FA): Measures the accuracy of the unlearned model $\theta_{u}$ on $\mathcal{D}_{f}$. Retain Accuracy (RA): Measures the accuracy of the unlearned model $\theta_{u}$ on $\mathcal{D}_{r}$. Test Forget Accuracy (TFA): Classification accuracy on the test forget set $\mathcal{D}^{\text{te}}_{f}$. Similar to FA, this evaluates the model’s ability to forget removed classes beyond the training data. Test Retain Accuracy (TRA): Classification accuracy on the test retain set $\mathcal{D}^{\text{te}}_{r}$. It indicates how effectively the model’s knowledge generalizes to unseen data of the remaining classes. Following [39], we introduce the Average GAP of Logit-based Metrics (AGL), which quantifies the discrepancy between the given model (e.g., the unlearned model $\theta_{u}$) and the golden standard $\theta_{r}$. This metric aggregates the logit-based evaluations:
$AGL=$ $\Bigl{(}1-\mathrm{G}(\theta_{u},\mathcal{D}_{f})\Bigr{)}\times\Bigl{(}1-\mathrm{G}(\theta_{u},\mathcal{D}_{r})\Bigr{)}\times\Bigl{(}1-\mathrm{G}(\theta_{u},\mathcal{D}^{\text{te}}_{f})\Bigr{)}\times\Bigl{(}1-\mathrm{G}(\theta_{u},\mathcal{D}^{\text{te}}_{r})\Bigr{)},$ (1)
@@
(a) Original ($\theta_{o}$)
(b) Retrained ($\theta_{r}$)
(c) FT
(d) GA [32]
(e) RL [13]
(f) SalUn [12]
Figure 3: We visualize the feature representations from the original ($\theta_{o}$) retrained ($\theta_{r}$), and unlearned models ($\theta_{u}$) on a subset of ImageNet-1K using ResNet-50. Unlike $\theta_{r}$, which serves as the gold standard, the forget classes (highlighted in red) in the existing unlearning baselines (e.g., (d), (e), and (f)) are severely dispersed throughout the entire class distribution.
@@
Representation-based evaluation metrics.   To overcome the limitations of logit-based evaluation, we propose an additional representation-based approach that assesses unlearning effectiveness by comparing features extracted from the encoder (e.g., after the average pooling layer in ResNet) between the unlearned and retrained models. This enables a fine-grained analysis of representational changes beyond what output-level metrics can capture. We use the following two metrics for the representation-based evaluation: First, we use Centered Kernel Alignment (CKA) [20] to measure the similarity between the feature representations of $\theta_{u}$ and $\theta_{r}$ as well as $\theta_{u}$ and $\theta_{o}$. Second, we measure $k$-Nearest Neighbors Accuracy ($k$-NN) [9] to assess the representational quality of $\theta_{u}$. We employ a $k$-NN classifier, which takes features of the test data from a model and performs classification based on the neighborhood features. Additionally, we also employ t-SNE [33] to visualize feature distributions. More details are provided in Sec. A.2 in the Appendix.
Model and datasets.   To ensure a more accurate and meaningful evaluation of unlearning algorithms considering real-world scenarios, we employed the ResNet-50 [18] model pre-trained on the ImageNet-1K [10] as the basic setting.
Unlearning baselines.   We evaluate several representative machine unlearning algorithms in the classification task, which can be categorized into two groups: 1) Unlearning without $\mathcal{D}_{r}$, including Gradient Ascent (GA)[32], Random Labeling (RL)[13], Pseudo Labeling (PL)[7], and SalUn[12]; and 2) unlearning with $\mathcal{D}_{r}$, including DUCK [11], Contrastive Unlearning (CU) [38], SCRUB [23], and SCAR [2]. Note that SalUn and SCAR are known as current state-of-the-art methods in each category. A more detailed explanation of each baseline can be found in Sec. A.1 in the Appendix.
Unlearning scenarios.   In our experiments and analysis, we consider two unlearning scenarios. The first is Random-$N$ Class-wise Forgetting, where we randomly select $N$ classes to remove. This scenario is a standard approach to evaluate unlearning algorithms. To provide a more comprehensive assessment, we introduce a second scenario, Top-$N$ Class-wise Forgetting, in which we specifically target classes that exhibit semantic similarity to the downstream task classes. The selection of $N$ classes is based on their relevance to the downstream task.
3.2 Motivation: Why Do We Need Representation-Based Metrics?
Current machine unlearning algorithms predominantly rely on logit-based evaluation metrics to assess their effectiveness. These algorithms aim to achieve classification accuracy similar to model trained only on retain set ($\theta_{r}$), and based on these metrics, they are considered to have successfully accomplished unlearning. However, this evaluation raises a critical research question: Do existing unlearning evaluations adequately reflect successful unlearning from a representation perspective?
t-SNE visualization.   We first conducted an analysis on several unlearning baselines using t-SNE in Fig. 3, which provides a visualization of feature embeddings from the encoder in a lower-dimensional space. While several unlearning algorithms (GA, RL, SalUn) achieve superior performance on logit-based evaluation metrics in small-scale settings, Fig. 3 reveals several important findings. First, (a) Original ($\theta_{o}$), (b) Retrain ($\theta_{r}$), and (c) Fine-Tuning (FT) models exhibit feature representations where each class forms distinct clusters. We expect that an effective unlearning algorithm should ideally produce feature representations similar to those of (b) ($\theta_{r}$). However, Figures (d)-(f) reveal that these approaches induce substantial collapse in feature representations, which deviate considerably from those of $\theta_{r}$. For all t-SNE visualization results, see Fig. 8 in the Appendix.
(a) CKA similarity of unlearning algorithms
(b) An analysis of $\theta_{u}^{last}$
Figure 4: Two analyses using CKA similarity are presented: (a) CKA similarity analysis of various unlearning algorithms. The x-axis shows the similarity to $\theta_{o}$ and the y-axis represents the similarity to $\theta_{r}$. In an ideal scenario, algorithms should be positioned near $\theta_{r}$. However, most algorithms exhibit a high similarity to $\theta_{o}$, suggesting that the transformation of representations during unlearning is suboptimal. (b) The X-axis depicts the feature similarity between fully unlearned model ($\theta_{u}$) and only the final layer unlearned model ($\theta_{u}^{last}$). Y-axis represents the gap in AGL scores between $\theta_{u}$ and $\theta_{u}^{last}$. The results indicate that existing algorithms primarily modify only the last layer while maintaining the original representation space.
CKA analysis.   In Fig. 4(a), we compare the Centered Kernel Alignment (CKA) similarity values between the unlearned ($\theta_{u}$) and retrained ($\theta_{r}$) models (on the y-axis) against those between the unlearned and original ($\theta_{o}$) models (on the x-axis), using ImageNet-1K. Note that CKA quantifies the representational similarity between different models, providing a measure of how similarly two models “see” or interpret the given data. Intuitively, if two models are aligned in their feature representations, they can be considered to share a similar perspective on the data. In the context of successful unlearning, we expect the CKA similarity between the unlearned and retrained models to be higher than that between the unlearned and original models. However, our findings reveal a critical issue: all unlearning baselines show a higher similarity to the original model’s representations than to the retrained model’s representations. In fact, certain baselines, such as GA, SalUn, and RL, exhibit very low similarity to both the original and retrained models.
Why does the unlearned model maintain high similarity to the original model?   Several unlearning baselines show high logit-based metric scores (i.e., minimal deviation from the retrained model) in our experimental setup, while still maintaining a strong resemblance to the original model’s representations. We hypothesize that this occurs because unlearning algorithms tend to focus on modifying the final layer to optimize logit-based metrics while leaving earlier representational layers largely unchanged. In Fig. 4(b), we compare the CKA similarities between unlearned models, where either all layers are tuned ($\theta_{u}$) or only the final layer is tuned ($\theta_{u}^{last}$), alongside the gap in logit-based AGL (Eq. 1) scores (y-axis). Ideally, the CKA similarity should be low (low x-values), which would correspond to a higher gap in AGL scores (high y-values). However, $\theta_{u}$ shows high CKA similarity to $\theta_{u}^{last}$ while presenting a relatively small gap in AGL, demonstrating a crucial limitation in the current evaluation approach.
In summary, the results from both t-SNE and CKA analyses highlight a key limitation in relying solely on logit-based metrics for unlearning evaluation. While some algorithms show high logit-based performance, their feature representations remain too similar to the original model, indicating insufficient unlearning. This underscores the need for representation-based metrics, which can more effectively capture the changes in feature representations and provide a more accurate measure of unlearning effectiveness, especially in large-scale, real-world applications.
4 A Unified Benchmark Framework for Machine Unlearning
To establish a rigorous and comprehensive evaluation of unlearning algorithms, we propose a unified framework that integrates both logit-based and representation-based metrics. This framework enables a more holistic assessment of unlearning performance by capturing both Feature Similarity and Transferability—two fundamental aspects of an unlearned model’s behavior. First, Feature Similarity quantifies the alignment of feature representations of the original and unlearned models on the same dataset. To measure this, we employ Centered Kernel Alignment (CKA), which provides a robust similarity measure between the learned representations of both models. Second, Transferability evaluates how well the unlearned model’s learned representations generalize to downstream tasks. We assess this by applying a $k$-Nearest Neighbors ($k$-NN) classifier to the encoder’s output features and assess performance on several downstream datasets. This allows us to measure how well features generalize beyond the original training distribution.
Table 1: The comparison of unlearning performance across different $\theta_{u}$ on ImageNet-1K under random 100 class-wise forgetting. The blue numbers represent the gap from $\theta_{r}$, demonstrating the efficacy of different unlearning strategies.
ImageNet-1K Office-Home CUB DomainNet-126 Office-Home CUB DomainNet-126 CKA($\theta_{u}$, $\theta_{r}$) $>$
Method FA RA TFA TRA Avg. Gap AGL $\uparrow$ $k$-NN CKA($\theta_{u}$, $\theta_{r}$) AGR $\uparrow$ H-LR $\uparrow$ CKA($\theta_{u}$, $\theta_{o}$)
Original 79.2 (79.2) 80.1 (4.1) 76.1 (76.1) 76.5 (0.9) 40.1 - 80.3 (3.0) 43.4 (6.5) 84.0 (2.0) 89.8 (10.2) 82.1 (17.9) 81.8 (18.2) - -
Retrained 0.0 (0.0) 76.0 (0.0) 0.0 (0.0) 75.6 (0.0) 0.0 1.00 77.3 (0.0) 36.9 (0.0) 82.0 (0.0) 100.0 (0.0) 100.0 (0.0) 100.0 (0.0) 1.00 1.00
FT 9.9 (9.9) 79.5 (3.5) 10.5 (10.5) 76.0 (0.4) 6.1 0.78 78.4 (1.1) 39.0 (2.1) 81.6 (0.4) 87.6 (12.4) 79.0 (21.0) 79.9 (20.1) 0.81 0.79
GA 1.5 (1.5) 11.4 (64.6) 1.5 (1.5) 11.3 (64.3) 33.0 0.12 29.7 (47.6) 10.4 (26.5) 42.2 (39.8) 8.3 (91.7) 9.8 (90.2) 10.2 (89.8) 0.06 0.08
RL 6.3 (6.3) 14.3 (62.7) 5.5 (5.5) 13.0 (62.6) 34.3 0.12 44.0 (33.3) 8.0 (28.9) 35.2 (46.8) 6.3 (93.7) 5.8 (94.2) 4.7 (95.3) 0.04 0.06
PL 1.0 (1.0) 79.5 (3.5) 1.0 (1.0) 76.9 (0.9) 1.6 0.94 80.3 (3.0) 43.3 (6.4) 84.0 (2.0) 91.6 (8.4) 84.7 (15.3) 84.5 (15.5) 0.84 0.89
SalUn 10.4 (10.4) 19.7 (56.3) 9.3 (9.3) 19.1 (56.5) 32.8 0.15 38.5 (38.8) 8.0 (28.9) 42.5 (39.5) 9.7 (90.3) 8.0 (92.0) 8.5 (91.5) 0.06 0.09
DUCK 0.9 (0.9) 74.6 (1.4) 0.9 (0.9) 74.5 (1.1) 1.1 0.96 79.8 (2.5) 39.0 (2.1) 82.5 (0.7) 90.7 (9.3) 83.2 (16.8) 84.9 (15.1) 0.85 0.90
CU 2.1 (2.1) 73.9 (2.1) 2.2 (2.2) 73.3 (2.3) 2.2 0.92 75.8 (1.5) 33.9 (3.0) 80.1 (1.9) 85.6 (14.4) 75.9 (24.1) 76.1 (23.9) 0.78 0.84
SCAR 4.2 (4.2) 80.5 (4.5) 3.9 (3.9) 77.4 (1.8) 3.6 0.87 78.0 (0.7) 42.8 (5.9) 83.1 (1.1) 74.2 (25.8) 65.4 (34.6) 58.8 (41.2) 0.64 0.74
SCRUB 1.1 (1.1) 67.3 (8.7) 1.1 (1.1) 65.7 (9.9) 5.2 0.80 74.7 (2.6) 42.2 (5.3) 80.9 (1.1) 68.9 (51.1) 63.8 (36.2) 52.8 (47.2) 0.60 0.69
Unified metric.   To establish a comprehensive evaluation of unlearning algorithms, we propose a unified measure that integrates both logit-based and representation-based evaluations. For representation-based evaluation, we use diverse downstream datasets, $\mathcal{D}_{\text{down}}$, which differ from the original train set distribution (e.g., ImageNet-1K). Inspired by transfer learning evaluation protocols [29], this approach enables us to assess both the quality and generalizability of learned representations. Specifically, we evaluate models on three downstream datasets: Office-Home [34], CUB [35], and DomainNet-126 [30]. More details of datasets are provided in Sec A.3 in the Appendix.
To formally quantify representation changes, we define two key metrics: First, $\mathrm{G_{kNN}}$ measures the difference in classification accuracy ($\text{A}^{k}$) of a $k$-NN classifier between $\theta_{u}$ and $\theta_{r}$ using a downstream dataset:
$\mathrm{G_{kNN}}(\theta_{u},\mathcal{D}_{\text{down}})=\bigl{|}\,\text{A}^{k}(\theta_{u},\mathcal{D}_{\text{down}})-\text{A}^{k}(\theta_{r},\mathcal{D}_{\text{down}})\bigr{|}.$
This metric evaluates the transfer learning ability of the given model’s encoder compared to that of the retrained model. In addition, we quantify the representational similarity between the given and retrained models for the given downstream dataset, by using $\mathrm{CKA}_{\text{AGR}}$: $\mathrm{CKA}_{\text{AGR}}(\theta_{u},\mathcal{D}_{\text{down}}),$ which measures the representational similarity between the given and retrained models over the downstream dataset.
Using these metrics, we define the Average GAP of Representation-based metrics (AGR) as:
@@ 4 A Unified Benchmark Framework for Machine Unlearning
$AGR=\Bigl{(}1-\mathrm{G_{kNN}}(\theta_{u},\mathcal{D}_{\text{down}})\Bigr{)}\times\mathrm{CKA}_{\text{AGR}}(\theta_{u},\mathcal{D}_{\text{down}}).$
This formulation balances the impact of both feature divergence and transferability degradation, providing a more nuanced measure of unlearning effectiveness from a representation perspective.
Finally, we introduce a unified evaluation measure, $H\text{-}LR$, which harmonizes both logit-based (AGL) and representation-based (AGR) metrics into a single, unified evaluation measure:
$H\text{-}LR=\frac{2}{\frac{1}{AGL}+\frac{1}{AGR}}.$
This harmonic mean formulation ensures that both classification performance and representational divergence are jointly considered, making $H\text{-}LR$ a robust and comprehensive metric for evaluating unlearning algorithms. Intuitively, a high H-LR score indicates that an unlearning algorithm achieves both output-level forgetting and meaningful representation-level divergence from the original model. In contrast, algorithms that modify only the classifier while leaving representations unchanged will score high on AGL but low on AGR, leading to a low H-LR. This unified metric penalizes such superficial unlearning.
4.1 Re-evaluation of Unlearning Baselines under Large-Scale Unlearning Scenarios
In this section, we present experimental results for random class-wise unlearning in large-scale settings, evaluating both logit-based and representation-based performance.
Small-scale vs. large-scale unlearning.   Several unlearning algorithms have demonstrated strong performance in small-scale settings, as evaluated by logit-based metrics. As shown in Fig. 5(a), both RL and SalUn achieve superior AGL on CIFAR-10. However, the figure also clearly highlights a significant performance degradation on ImageNet-1K, underscoring the challenges associated with scaling unlearning algorithms to larger datasets. Fig. 5(b) further presents both AGL and AGR scores in large-scale unlearning experiments. AGL scores, measured using ImageNet-1K, are accompanied by averaged AGR scores, which are calculated using multiple downstream datasets, $\mathcal{D}_{\text{down}}$. This figure emphasizes the limitations of both existing unlearning algorithms and the evaluation methods currently in use. First, three algorithms—GA, RL, and SalUn—demonstrate a significant drop in performance in the large-scale unlearning scenario. Second, all baseline algorithms consistently exhibit low AGR compared to AGL, indicating their failure to effectively unlearn from a representational perspective.
(a) Large vs. small scale
(b) AGL vs. AGR
Figure 5: (a) AGL results for unlearning experiments on pretrained ResNet-50 models using CIFAR-10 and ImageNet-1K. We conduct unlearning on a pretrained model for one randomly selected class from CIFAR-10 and 100 randomly selected classes from ImageNet-1K. (b) AGL and AGR results for unlearning experiments on ImageNet-1K pretrained ResNet-50 models. We conduct unlearning on a pretrained model for 100 randomly selected classes from ImageNet-1K and report the corresponding AGL and AGR values. Notably, following SalUn’s protocol, we excluded test set accuracy when calculating AGL in both (a) and (b)
Experimental results in large-scale unlearning.   Table 1 presents the experimental results for all baselines and evaluation methods in our large-scale unlearning scenarios. Through logit-based evaluation and AGL scores, we observe that PL (which does not use $\mathcal{D}_{r}$) achieves performance comparable to or even surpassing state-of-the-art algorithms that do utilize $\mathcal{D}_{r}$ (i.e., DUCK, CU, SCAR, and SCRUB). This result marks a notable departure from the trends observed in small-scale unlearning experiments [11, 26, 2, 23]. In representation-based evaluation (measured by AGR), PL consistently outperforms other algorithms, achieving superior H-LR among the baselines. These results challenge prior expectations, showing that many state-of-the-art algorithms, such as SalUn, SCAR, and SCRUB, degrade in large-scale setup, performing worse from both logit and representation perspectives.
4.2 Experiments on Top Class-wise Forgetting
Why top class-wise forgetting is necessary?   In the random-class forgetting scenario, the CKA similarity between $\theta_{o}$ and $\theta_{r}$ reaches a high value of 88%, as shown in Fig. 4(a). We hypothesize that this high similarity arises from the presence of multiple alternative classes within ImageNet, which may preserve similar representational knowledge. Consequently, this results in a significant overlap between both models, presenting challenges in verifying the effectiveness of unlearning from a representational perspective.
To address this, we propose a novel top class-wise forgetting approach, where forgetting classes are selected based on semantic similarity criteria. This ensures that no alternative classes with similar representations remain in $\mathcal{D}_{r}$, leading to a substantial difference between $\theta_{r}$ and $\theta_{o}$. In other words, this setup makes unlearning more challenging, allowing for a more rigorous evaluation of unlearning algorithms. As shown in Fig. 9 in the Appendix, this approach minimizes CKA similarity between $\theta_{o}$ and $\theta_{r}$, thereby facilitating a more reliable and straightforward evaluation of unlearning.
Similarity-based Top-class Selection. In selecting semantically similar classes, we define $\mathcal{D}_{f}$ as the classes most similar to those in downstream datasets. To measure similarity, we extract feature maps for each class from the average pooling layer of ResNet-50 and compute cosine similarity between the class-wise averaged feature maps. We rank the classes and select the top 100 most similar ones for $\mathcal{D}_{f}$. For example, in the case of Office-Home, which consists of 65 classes, we identify the most similar ImageNet-1K classes to this dataset as the forget set. We extend the next 35 most similar classes, creating a 100-class forget set. If unlearning is successful, we expect to observe corresponding changes in performance on related classes in downstream tasks.
Random vs. Top class-wise forgetting.   We conduct extensive experiments comparing random and top class-wise unlearning across several datasets. The results are presented in Fig. 6. An ideal unlearning should show minimal gaps in both AGL and AGR, regardless of unlearning scenarios. As shown in Fig. 6(a), PL consistently achieves high and comparable AGL scores across all settings. However, its AGR scores exhibit significant variation depending on the unlearning scenario, highlighting its weakness in top class-wise unlearning. Furthermore, both SalUn (Fig. 6(b)) and CU (Fig. 6(d)) show considerably degraded performance in terms of both AGL and AGR in most top class-wise unlearning cases. Additionally, even though DUCK (Fig. 6(c)) achieves relatively stable performance compared to other baselines, we observe that most unlearning methods exhibit significantly lower AGR scores compared to the corresponding AGL scores, especially in the CUB Top case.
(a) PL
(b) SalUn
(c) DUCK
(d) CU
Figure 6: AGL and AGR comparison on unlearning baselines, using Random (R), Office-Home (OH), CUB (C), and DomainNet-126 (D) to select 100 similar classes for top class-wise unlearning.
(a) AGL scores
(b) AGR scores
Figure 7: AGL and AGR scores in Random and Top class-wise forgetting scenarios on PL. (a) AGL score archives steadily high in all settings. (b) AGR scores decrease, as the number of forgetting classes increases. This is evident in top class-wise forgetting, where $\theta_{r}$ has a low similarity to $\theta_{o}$.
Need for considering both AGL and AGR, respectively.   Fig. 6 reveals that unlearning algorithms exhibit unexpected behaviors, which become evident when considering both AGR and AGL scores. Despite PL and DUCK often showing little disparity between these scores, there is an inherent need to assess these measures independently to fully understand the performance of the unlearning algorithms. To demonstrate this, we extended the evaluation of PL by conducting experiments that unlearn 100 to 300 top semantically similar classes. As the number of forgetting classes increases, particularly in top class-wise forgetting, the nearest classes diverge significantly from the original ones. Fig. 7 illustrates that AGL scores remain largely unaffected by the number of forgetting classes. However, AGR scores notably decrease as the number of forgotten classes increases, indicating that PL still retains substantial similarity to $\theta_{o}$ ’s representation. These findings not only highlight the discrepancy between the unlearning performed by these algorithms and the expected results but also emphasize the need for our proposed multi-faceted evaluation approach. Fig. 10 in the appendix shows similar results on DUCK.
More experimental results are reported in the Appendix, including evaluations of alternative models (e.g., Swin-T [24]) under our experimental setup.
5 Concluding Remarks
Main contribution.   We propose a unified benchmark for evaluating machine unlearning under realistic, large-scale conditions. Our framework incorporates representation-based metrics (CKA, $k$-NN) and a novel Top Class-wise Forgetting scenario to reveal the limitations of current methods beyond traditional logit-based metrics. We show that many algorithms fail to significantly alter internal representations, despite showing high output-level scores. Our benchmark offers a scalable and reproducible evaluation protocol, laying the groundwork for more rigorous and reliable unlearning research. All code and datasets are publicly released to support community adoption.
Limitations.   Our benchmark is limited to classification tasks with accessible encoders and is not directly applicable to generative or black-box models. CKA and $k$-NN metrics, while informative, may not capture all semantic shifts. Future extensions should address dynamic unlearning and broader task types.
Broader impact.   Our benchmark enables more accountable and rigorous evaluation of unlearning algorithms, critical for privacy-sensitive applications such as healthcare or personalized AI. However, there’s a risk that future methods might overfit to benchmark metrics without achieving true forgetting. We advocate combining our benchmark with formal privacy guarantees and adversarial robustness checks to ensure safe deployment.
Appendix A Detail Descriptions
A.1 Unlearning Baselines

### v3

@@ 3.2 Logit-based Evaluation Metrics
As a classical evaluation protocol, unlearning models are evaluated based on the following metrics: Forget Accuracy (FA): Measures the accuracy of the unlearned model $\theta_{u}$ on $\mathcal{D}_{f}$. Retain Accuracy (RA): Measures the accuracy of the unlearned model $\theta_{u}$ on $\mathcal{D}_{r}$. Test Forget Accuracy (TFA): Classification accuracy on the test forget set $\mathcal{D}^{\text{te}}_{f}$. Similar to FA, this evaluates the model’s ability to forget removed classes beyond the training data. Test Retain Accuracy (TRA): Classification accuracy on the test retain set $\mathcal{D}^{\text{te}}_{r}$. It indicates how effectively the model’s knowledge generalizes to unseen data from the remaining classes. Following [53], we introduce the Average GAP of Logit-based Metrics (AGL), which quantifies the discrepancy between the given model (e.g., the unlearned model $\theta_{u}$) and the gold standard $\theta_{r}$. This metric aggregates the logit-based evaluations:
$AGL=$ $\Bigl(1-\mathrm{G}(\theta_{u},\mathcal{D}_{f})\Bigr)\times\Bigl(1-\mathrm{G}(\theta_{u},\mathcal{D}_{r})\Bigr)$ (1)
$\times\Bigl(1-\mathrm{G}(\theta_{u},\mathcal{D}^{\text{te}}_{f})\Bigr)\times\Bigl(1-\mathrm{G}(\theta_{u},\mathcal{D}^{\text{te}}_{r})\Bigr),$
@@
Membership Inference Attack (MIA). Membership Inference Attack efficacy is measured using a prediction confidence-based attack method [42]. The process includes two phases: (1) training, where a balanced dataset from $\mathcal{D}_{r}$ and unseen test data is used to train the Support Vector Machine (SVM)-based MIA predictor; and (2) testing, where the trained MIA model evaluates the unlearning performance. It measures the unlearning effectiveness by applying the MIA binary classifier to the unlearned model $\theta_{u}$ on the forget set $\mathcal{D}_{f}$. The more the model predicts the forget set as test data, the higher the resulting MIA-efficacy. MIA-Efficacy evaluates how effectively unlearning removes information by using privacy attack methods. We report the MIA results in Table 5.
3.3 Representation-based Evaluation Metrics
@@
To overcome the limitations of logit-based evaluation, we propose an additional representation-based approach that assesses unlearning effectiveness by comparing features extracted from the encoder (e.g., after the average pooling layer in ResNet) between the unlearned and retrained models. This enables a fine-grained analysis of representational changes beyond what output-level metrics can capture. We use the following two metrics for the representation-based evaluation: First, we use Centered Kernel Alignment (CKA) [26] to measure the similarity between the feature representations of $\theta_{u}$ and $\theta_{r}$, as well as $\theta_{u}$ and $\theta_{o}$. Second, we measure $k$-Nearest Neighbors Accuracy ($k$-NN) [13] to assess the representational quality of $\theta_{u}$. We employ a $k$-NN classifier, which takes features of data from a targeted model and performs classification based on the neighborhood features. Additionally, we also employ t-SNE [45] to visualize feature distributions.
where $K$ and $L$ represent the Gram matrices of the two feature representations, and HSIC denotes the Hilbert-Schmidt independence criterion. By applying CKA, we measure the similarity between the feature representations of $\theta_{u}$ (the unlearned model), $\theta_{o}$ (the original model), and $\theta_{r}$ (the retrained model) across various downstream datasets. Note that higher CKA values indicate a greater degree of alignment, suggesting that the unlearned model retains the desired feature representations after unlearning.
$k$-Nearest Neighbors ($k$-NN). To evaluate the transferability of the unlearned model, we employ the $k$-NN method. Feature embeddings are extracted from the unlearned model, $\theta_{u}$, by feeding it the downstream data. These embeddings are then split, with 80% used to train a $k$-NN classifier and the remaining 20% used for testing. For each test sample, the classifier assigns a label based on the majority vote of its $k$ nearest neighbors from the training set. The $k$-NN Accuracy is computed as the proportion of correctly classified test samples. This evaluation provides insights into whether $\theta_{u}$ preserves meaningful, transferable feature embeddings that are effective for downstream tasks. We set $k=5$ in the $k$-Nearest Neighbors ($k$-NN) classifier in experiments, using standard cosine distance. Ties in the majority vote are broken by selecting the class with the lowest numerical index. However, we consider its impact negligible, as such occurrences were rare in our evaluation.
t-SNE visualization. To understand how unlearning affects a model’s internal features, we visualized the features using t-SNE. Instead of visualizing the full dataset, we selected a consistent, randomly chosen subset of classes from both the forget set and the retain set due to computational costs. This subset is identical across all unlearning algorithms, ensuring fair comparison. The t-SNE visualization of all unlearning baselines is shown in Fig. 12.
Average GAP of Representation-based metrics (AGR). AGR is designed to holistically assess representation-level unlearning by measuring two key aspects: the feature representational alignment with the retrained model, measured by $\mathrm{CKA}_{\text{AGR}}$, and the preservation of transferable knowledge, evaluated by $\mathrm{G_{kNN}}$. Details on these metrics ($\mathrm{CKA}_{\text{AGR}}$, $\mathrm{G_{kNN}}$) are introduced in Sec. 4.1. The AGR score is computed by combining these two components as follows:
@@
$AGR=\Bigl(1-\mathrm{G_{kNN}}(\theta_{u},\mathcal{D}_{\text{down}})\Bigr)\times\mathrm{CKA}_{\text{AGR}}(\theta_{u},\mathcal{D}_{\text{down}}).$ (3)
The calculation of these components depends on the unlearning scenario. For the Random Class-wise Forgetting scenario, we measure $\mathrm{G_{kNN}}$ and $\mathrm{CKA}_{\text{AGR}}$ across all downstream datasets and use their average values. However, for the Top Class-wise Forgetting scenario, the metrics are measured only on the semantically related downstream dataset. For instance, when evaluating a model under the CUB Top-100 Class-wise Forgetting scenario, we calculate $\mathrm{G_{kNN}}$ and $\mathrm{CKA}_{\text{AGR}}$ using only the CUB dataset.
3.4 Models and Datasets
Models and datasets. To ensure a more meaningful evaluation of unlearning algorithms considering real-world scenarios, we employ the ResNet-50 [21] model pre-trained on ImageNet-1K [14] as the large-scale setting. We further extend our experiments to include the Swin-T [31] and ConvNeXt [32] architectures.
Downstream datasets. For our evaluation framework, we employ three diverse downstream datasets: Office-Home [46]: A domain adaptation benchmark comprising distinct domains. We use Real World, Art, Clipart, and Product domains. It includes 65 classes with over 15,500 images, providing a robust basis for cross-domain transfer learning. CUB-200-2011 (CUB) [47]: The Caltech-UCSD Birds dataset features fine-grained classification with 200 bird species and 11,788 images, suitable for evaluating detailed representational quality. DomainNet-126 [38]: A multi-source domain dataset containing several domains. We use Clipart, Painting, Real, and Sketch domains. This dataset includes 126 classes.
3.5 Unlearning Baselines